# references.bib

### v3

@@ 3. Tautological classes
Then, for any line bundle $\mathcal{L}$ over $V/G$, we have an operation $\pi_{*}\mathbf{Q}_{V/L}\rightarrow\pi_{*}\mathbf{Q}_{V/L}$ of multiplication by the first Chern class $c_{1}(\mathcal{L})$. More precisely, we have the following lemma.
@@
We have a morphism of graded rings
[…]
@@
We have a canonical adjunction morphism
@@
Let $\mathfrak{g}$ be a reductive Lie algebra, $\mathfrak{h}\subseteq\mathfrak{g}$ an ideal and $\mathfrak{t}$ a Cartan subalgebra. Then, $\mathfrak{h}\cap\mathfrak{t}$ is a Cartan subalgebra of $\mathfrak{h}$.
@@
Let $\pi\colon\mathfrak{g}\rightarrow\mathfrak{g}/\mathfrak{h}$ be the projection. Since $\mathfrak{g}$ is reductive, we have a (noncanonical) decomposition $\mathfrak{g}\cong\mathfrak{h}\oplus(\mathfrak{g}/\mathfrak{h})$. If $\mathfrak{t}^{\prime}$ is a maximal torus of $\mathfrak{h}$ and $\mathfrak{t}^{\prime\prime}$ is a maximal torus of $\mathfrak{g}/\mathfrak{h}$, then $\mathfrak{t}^{\prime}\oplus\mathfrak{t}^{\prime\prime}$ is a maximal torus of $\mathfrak{g}$. Moreover, $\pi(\mathfrak{t})\cong\mathfrak{t}/\mathfrak{t}\cap\mathfrak{h}$ is an Abelian subalgebra of $\mathfrak{g}/\mathfrak{h}$ and $\mathfrak{t}\cap\mathfrak{h}$ is an Abelian subalgebra of $\mathfrak{h}$. By maximality of $\dim\mathfrak{t}^{\prime}$, we have
$\dim\pi(\mathfrak{t})=\dim\mathfrak{t}-\dim(\mathfrak{t}\cap\mathfrak{h})\geqslant\dim\mathfrak{t}-\dim\mathfrak{t}^{\prime}=\dim\mathfrak{t}^{\prime\prime}$
@@
It is necessary to consider the neutral component of $T\cap H$ as shown by the elementary case of $H=\{\pm 1\}\subseteq\mathrm{SL}_{2}(\mathbf{C})$, and where $T\subseteq\mathrm{SL}_{2}(\mathbf{C})$ is the standard torus.
@@
By choosing a decomposition $\mathrm{Lie}(G)\cong\mathrm{Lie}(H)\oplus\mathrm{Lie}(G/H)$, we obtain a decomposition $\mathrm{H}^{*}(\mathrm{pt}/G)\cong\mathrm{H}^{*}(\mathrm{pt}/(G/H))\otimes\mathrm{H}^{*}(\mathrm{pt}/H)$. At the sheaf level, the action of the lemma is obtained by pre-composing the morphism $a$ of Lemma 3.1 with the inclusion $\mathrm{H}^{*}(\mathrm{pt}/H)\rightarrow\mathrm{H}^{*}(\mathrm{pt}/G)$. ∎
@@
The map $\mathrm{H}^{*}(\mathrm{pt}/(G/H))\rightarrow\mathrm{H}^{*}(\mathrm{pt}/G)$ is the pullback in cohomology for the morphism of stacks $\mathrm{pt}/(G/H)\rightarrow\mathrm{pt}/G$ induced by the surjective morphism $G\rightarrow G/H$. ∎
@@
Let $G$ be an algebraic group, and $V$ a representation of $G$. Let $H$ be a normal subgroup of $G$ acting trivially on $V$. Then, the morphism
@@
We may assume that $V=\mathrm{pt}$, that $G$ is connected and by quotienting out by the unipotent radical of $G$, that $G$ is reductive. We let $\mathfrak{g}=\mathrm{Lie}(G)$ and $\mathfrak{h}=\mathrm{Lie}(H)$. We let $\mathfrak{t}\subseteq\mathfrak{g}$ be a Cartan subalgebra. We have $\mathfrak{g}\cong\mathfrak{h}\oplus(\mathfrak{g}/\mathfrak{h})$ and therefore a decomposition of the Weyl group $W_{\mathfrak{g}}\cong W_{\mathfrak{h}}\times W_{\mathfrak{g}/\mathfrak{h}}$. We let $\pi\colon\mathfrak{g}\rightarrow\mathfrak{g}/\mathfrak{h}$ be the projection. The morphism of the proposition can then be identified with the isomorphism
$\mathrm{Sym}((\mathfrak{t}\cap\mathfrak{h})^{*})^{W_{\mathfrak{h}}}\otimes\mathrm{Sym}(\pi(\mathfrak{t})^{*})^{W_{\mathfrak{g}/\mathfrak{h}}}\rightarrow\mathrm{Sym}(\mathfrak{t}^{*})^{W}.$
4. Cohomological integrality for symmetric representations of reductive groups
We fix a symmetric representation $V$ of a connected reductive group $G$.
@@ 4.1. Weyl groups
Let $\lambda\colon\mathbf{G}_{\mathrm{m}}\rightarrow T$. We let $G^{\lambda}\subseteq G$ be the corresponding Levi subgroup and $G_{\lambda}$ the intersection of the kernel of the action of $G^{\lambda}$ on $V^{\lambda}$ and the center of $G^{\lambda}$. We let $W^{\lambda}=N_{G^{\lambda}}(T)$ be the Weyl group of $G^{\lambda}$ and $W_{\lambda}\coloneqq\{w\in W\mid V^{w\cdot\lambda}=V^{\lambda}\text{ and }G^{w\cdot\lambda}=G^{\lambda}\}=\{w\in W\mid w\cdot\overline{\lambda}=\overline{\lambda}\}$.
@@
$W_{\lambda}$ is a subgroup of $W$.
@@
We let $\dot{w}\in N_{G}(T)$ be a lift of $w\in W$. Then, $V^{w\cdot\lambda}=\dot{w}V^{\lambda}$. The statement of the lemma follows then immediately from the fact that if $w,w^{\prime}\in W$, and $\dot{w},\dot{w^{\prime}}\in N_{G}(T)$ are respective lifts, then, $\dot{w}\dot{w^{\prime}}$ is a lift of $ww^{\prime}$ and $\dot{w}^{-1}$ a lift of $w^{-1}$. ∎
@@
We have natural inclusions $W^{\lambda}\subseteq W_{\lambda}\subseteq W$.
@@
The inclusion $W_{\lambda}\subseteq W$ holds by definition of $W_{\lambda}$ (see Lemma 4.1). The inclusion $W^{\lambda}\subseteq W$ comes from the inclusion $N_{G^{\lambda}}(T)\subseteq N_{G}(T)$ since $G$ and $G^{\lambda}$ share the same maximal torus $T$.
Let $w\in W^{\lambda}$ and $\dot{w}\in N_{G^{\lambda}}(T)$ a lift of $w$. Then, $\dot{w}G^{\lambda}\dot{w}^{-1}=G^{w\cdot\lambda}=G^{\lambda}$ and $\dot{w}V^{\lambda}=V^{w\cdot\lambda}=V^{\lambda}$ and so $w\in W_{\lambda}$. ∎
Lemma 4.3.
The group $W_{\lambda}$ normalises the group $W^{\lambda}$ inside $W$.
Proof.
It suffices to check that if $w\in W_{\lambda}$ and $w^{\prime}\in W^{\lambda}$, and $\dot{w}\in N_{G}(T)$, $\dot{w^{\prime}}\in N_{G^{\lambda}}(T)$ are any lifts, then $\dot{w}\dot{w^{\prime}}\dot{w}^{-1}\in G^{\lambda}$. This comes from the fact that $\dot{w}$ and $\lambda$ commute and $\lambda$ fixes $G^{\lambda}$ and so $\dot{w}\dot{w^{\prime}}\dot{w}^{-1}$ is fixed by $\lambda$. ∎
@@
As the Weyl group $W$ of $G$ acts on $\mathrm{H}^{*}(\mathrm{pt}/T)$. We just have to check that the induced action of the subgroup $W_{\lambda}$ preserves $W^{\lambda}$-invariants. This is a consequence of Lemma 4.3. ∎
@@
The group $W_{\lambda}$ acts naturally on $\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$.
@@
By definition, $G_{\lambda}$ is preserved by $W_{\lambda}$. ∎
We let $\overline{G^{\lambda}}\coloneqq G^{\lambda}/G_{\lambda}$. Since $G_{\lambda}$ is a normal subgroup of $G^{\lambda}$, $\overline{G^{\lambda}}$ is a reductive group. Since $G_{\lambda}$ is contained in the center of $G^{\lambda}$, the Weyl group of $\overline{G^{\lambda}}$ is isomorphic to $W$.
@@
The group $W_{\lambda}$ acts naturally on $\mathrm{H}^{*}(V^{\lambda}/\overline{G^{\lambda}})$.
@@
We have $\mathrm{H}^{*}(V^{\lambda}/\overline{G^{\lambda}})\cong\mathrm{H}^{*}(\mathrm{pt}/(T/G_{\lambda}))^{W^{\lambda}}$.
The restriction to $W_{\lambda}$ of the $W$ action on $T$ preserves $G_{\lambda}$ since $G_{\lambda}$ is in the center of $G^{\lambda}$. Therefore, we obtain a $W_{\lambda}$-action on $T/G_{\lambda}$. It induces a $W_{\lambda}$-action on $\mathrm{H}^{*}(\mathrm{pt}/(T/G_{\lambda}))$. The fact that $W_{\lambda}$ normalises $W^{\lambda}$ (Lemma 4.3) implies that $\mathrm{H}^{*}(\mathrm{pt}/(T/G_{\lambda}))^{W^{\lambda}}$ is preserved by the $W_{\lambda}$-action. ∎
By combining Corollary 4.4, Lemma 4.5 and Lemma 4.6, we obtain the following lemma.
Lemma 4.7.
The $W_{\lambda}$-action on $\mathrm{H}^{*}(\mathrm{pt}/G^{\lambda})\cong\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})$ obtained in Corollary 4.4 and the $W_{\lambda}$-action on $\mathrm{H}^{*}(V^{\lambda}/\overline{G^{\lambda}})\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ obtained by the tensor product of the actions given in Lemmas 4.5 and 4.6 coincide via the isomorphism given in Proposition 3.7.∎
@@
There exists a character $\varepsilon_{V,\lambda}\colon W_{\lambda}\rightarrow\{\pm 1\}$ such that for any $w\in W_{\lambda}$, $w(k_{\lambda})=\varepsilon_{V,\lambda}(w)k_{\lambda}$.
@@
It suffices to prove the existence and unicity of the sign $\varepsilon_{V,\lambda}(w)\in\{\pm 1\}$ satisfying the equality of the proposition, for any $w\in W_{\lambda}$. The multiplicativity then follows from the unicity of this sign.
The set of weights $\mathcal{W}^{\lambda=0}(V)=\mathcal{W}(V^{\lambda})$ only depends on $\overline{\lambda}$ and so is stable under the action of $W_{\lambda}$. Therefore, $\mathcal{W}^{\lambda>0}(V)\sqcup\mathcal{W}^{\lambda<0}(V)=\mathcal{W}(V)\smallsetminus\mathcal{W}^{\lambda=0}(V)$ is stable under $W_{\lambda}$. Also, by symmetry, $\mathcal{W}^{\lambda>0}(V)=-\mathcal{W}^{\lambda<0}(V)$. Therefore,
$\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V)\sqcup\mathcal{W}^{\lambda<0}(V)}\alpha=(-1)^{\#\mathcal{W}^{\lambda>0}(V)}\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V)}\alpha^{2}.$
By applying this equality to $\mathfrak{g}$ under the adjoint action, seen as a symmetric representation of $G$, we obtain
$\prod_{\alpha\in\mathcal{W}^{\lambda>0}(\mathfrak{g})\sqcup\mathcal{W}^{\lambda<0}(\mathfrak{g})}\alpha=(-1)^{\#\mathcal{W}^{\lambda>0}(\mathfrak{g})}\prod_{\alpha\in\mathcal{W}^{\lambda>0}(\mathfrak{g})}\alpha^{2}.$
Therefore,
$(-1)^{\#\mathcal{W}^{\lambda>0}(V)-\#\mathcal{W}^{\lambda>0}(\mathfrak{g})}k_{\lambda}^{2}=\frac{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V)\sqcup\mathcal{W}^{\lambda<0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(\mathfrak{g})\sqcup\mathcal{W}^{\lambda<0}(\mathfrak{g})}\alpha}.$
By applying $w\in W_{\lambda}$, we obtain $w(k_{\lambda}^{2})=k_{\lambda}^{2}$. Therefore, $w(k_{\lambda})=\pm k_{\lambda}$, which proves the existence of the sign. The unicity is straighforward. ∎
@@
Let $\lambda\in X_{*}(T)$, $w\in W_{\lambda}$ and $f\in\mathcal{H}_{\lambda}$. Then, $\mathrm{Ind}_{\lambda}(w\cdot f)=\varepsilon(w)\mathrm{Ind}_{\lambda}(f)$.
@@
We now give an elementary result from the representation theory of finite groups.
@@
There are essentially three things to check:
  1. (1)
    $p_{\chi}$ is an involution: $p_{\chi}\circ p_{\chi}=p_{\chi}$,
  2. (2)
    $p_{\chi}(f)=f$ for $f\in V^{\chi}$,
  3. (3)
    $\mathrm{im}(p_{\chi})\subseteq V^{\chi}$,
which are routine. ∎

### v6

@@ 3. Action of tautological classes on the cohomology of the stack
Then, for any line bundle $\mathcal{L}$ over $V/G$, we have an operation $\pi_{*}\mathbf{Q}_{V/G}\rightarrow\pi_{*}\mathbf{Q}_{V/G}[-2]$ of multiplication by the first Chern class $c_{1}(\mathcal{L})$. More precisely, we have the following lemma.
@@
There is a morphism of graded rings
$\mathrm{H}^{*}(\mathrm{pt}/G)\xrightarrow{a}\mathrm{RHom}(\pi_{*}\mathbf{Q}_{V/G},\pi_{*}\mathbf{Q}_{V/G})$
@@
We have a canonical adjunction isomorphism
@@
Let $\mathfrak{g}$ be a reductive Lie algebra, $\mathfrak{h}\subseteq\mathfrak{g}$ a Lie ideal and $\mathfrak{t}$ a Cartan subalgebra. Then, $\mathfrak{h}\cap\mathfrak{t}$ is a Cartan subalgebra of $\mathfrak{h}$.
@@
Let $\pi\colon\mathfrak{g}\rightarrow\mathfrak{g}/\mathfrak{h}$ be the projection. Since $\mathfrak{g}$ is reductive, we have a (noncanonical) decomposition $\mathfrak{g}\cong\mathfrak{h}\oplus(\mathfrak{g}/\mathfrak{h})$ as Lie algebras. If $\mathfrak{t}^{\prime}$ is a Cartan subalgebra of $\mathfrak{h}$ and $\mathfrak{t}^{\prime\prime}$ is a Cartan subalgebra of $\mathfrak{g}/\mathfrak{h}$, then $\mathfrak{t}^{\prime}\oplus\mathfrak{t}^{\prime\prime}$ is a Cartan subalgebra of $\mathfrak{g}$. Moreover, $\pi(\mathfrak{t})\cong\mathfrak{t}/\mathfrak{t}\cap\mathfrak{h}$ is an Abelian subalgebra of $\mathfrak{g}/\mathfrak{h}$ and $\mathfrak{t}\cap\mathfrak{h}$ is an Abelian subalgebra of $\mathfrak{h}$. By maximality of $\dim\mathfrak{t}^{\prime}$, we have
$\dim\pi(\mathfrak{t})=\dim\mathfrak{t}-\dim(\mathfrak{t}\cap\mathfrak{h})\geqslant\dim\mathfrak{t}-\dim\mathfrak{t}^{\prime}=\dim\mathfrak{t}^{\prime\prime}\,.$
@@
In the group situation, it is necessary to consider the neutral component of $T\cap H$ as shown by the elementary case of $H=\{\pm 1\}\subseteq\mathrm{SL}_{2}(\mathbf{C})$, and where $T\subseteq\mathrm{SL}_{2}(\mathbf{C})$ is the standard torus of diagonal matrices.
@@
By choosing a decomposition $\mathrm{Lie}(G)\cong\mathrm{Lie}(H)\oplus\mathrm{Lie}(G/H)$, we obtain a decomposition $\mathrm{H}^{*}(\mathrm{pt}/G)\cong\mathrm{H}^{*}(\mathrm{pt}/(G/H))\otimes\mathrm{H}^{*}(\mathrm{pt}/H)$, giving an action of $\mathrm{H}^{*}(\mathrm{pt}/H)$ on $\mathrm{H}^{*}(\mathrm{pt}/G)$ by cup-product on the second summand. At the sheaf level, the action of the lemma is obtained by pre-composing the morphism $a$ of Lemma 3.1 with the inclusion $\mathrm{H}^{*}(\mathrm{pt}/H)\rightarrow\mathrm{H}^{*}(\mathrm{pt}/G)$. ∎
@@
The map $\mathrm{H}^{*}(\mathrm{pt}/(G/H))\rightarrow\mathrm{H}^{*}(\mathrm{pt}/G)$ is the pullback in cohomology for the morphism of stacks $\mathrm{pt}/G\rightarrow\mathrm{pt}/(G/H)$ induced by the surjective morphism of groups $G\rightarrow G/H$. ∎
@@
Let $G$ be a reductive algebraic group, and $V$ a finite-dimensional representation of $G$. Let $H$ be a normal subgroup of $G$ acting trivially on $V$. Then, the morphism
@@
We let $\mathfrak{g}=\mathrm{Lie}(G)$ and $\mathfrak{h}=\mathrm{Lie}(H)$. We let $\mathfrak{t}\subseteq\mathfrak{g}$ be a Cartan subalgebra. We have $\mathfrak{g}\cong\mathfrak{h}\oplus(\mathfrak{g}/\mathfrak{h})$ and therefore a decomposition of the Weyl group $W_{\mathfrak{g}}\cong W_{\mathfrak{h}}\times W_{\mathfrak{g}/\mathfrak{h}}$. We let $\pi\colon\mathfrak{g}\rightarrow\mathfrak{g}/\mathfrak{h}$ be the projection. The morphism of the proposition can then be identified with the isomorphism
$\mathrm{Sym}((\mathfrak{t}\cap\mathfrak{h})^{*})^{W_{\mathfrak{h}}}\otimes\mathrm{Sym}(\pi(\mathfrak{t})^{*})^{W_{\mathfrak{g}/\mathfrak{h}}}\rightarrow\mathrm{Sym}(\mathfrak{t}^{*})^{W}\,$,
see also Lemma 3.2. ∎
4. Cohomological integrality for weakly symmetric representations of reductive groups
We fix a weakly symmetric representation $V$ (Definition 2.3) of a connected reductive group $G$ and a maximal torus $T\subseteq G$.
@@ 4.1. Weyl groups
Let $\lambda\colon\mathbf{G}_{\mathrm{m}}\rightarrow T$. We let $G^{\lambda}\subseteq G$ be the corresponding Levi subgroup (§2.1) and $G_{\lambda}\coloneqq Z(G^{\lambda})\cap\ker(G^{\lambda}\rightarrow\mathrm{GL}(V^{\lambda}))$ the intersection of the kernel of the action of $G^{\lambda}$ on $V^{\lambda}$ and the center of $G^{\lambda}$. We let $W^{\lambda}=N_{G^{\lambda}}(T)$ be the Weyl group of $G^{\lambda}$ and $W_{\lambda}\coloneqq\{w\in W\mid V^{w\cdot\lambda}=V^{\lambda}\text{ and }G^{w\cdot\lambda}=G^{\lambda}\}=\{w\in W\mid w\cdot\overline{\lambda}=\overline{\lambda}\}$. The group $W_{\lambda}$ may be thought of as a kind of relative Weyl group.
@@
Let $G_{\lambda}^{\circ}$ be the neutral component of $G_{\lambda}$. Let $k=\dim G_{\lambda}$. Then, $G_{\lambda}^{\circ}\cong(\mathbf{C}^{*})^{k}$ for some $k\geqslant 0$ and the natural pullback map $\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\rightarrow\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda}^{\circ})$ is an isomorphism.
@@
The group $G_{\lambda}$ is a subgroup of the center of $G^{\lambda}$, and so it is a diagonalisable algebraic group. Therefore, its neutral component is a torus. We have a splitting $G\cong G_{\lambda}^{\circ}\times\pi_{0}(G_{\lambda})$ [Hum12, Theorem p.104] and so, by the Künneth formula,
$\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\cong\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda}^{\circ})\otimes\mathrm{H}^{*}(\mathrm{pt}/\pi_{0}(G_{\lambda}))\,.$
@@
The group $W_{\lambda}$ is a subgroup of $W$ and it contains $W^{\lambda}$ as a normal subgroup.
@@
We let $\dot{w}\in N_{G}(T)$ be a lift of $w\in W$. Then, $V^{w\cdot\lambda}=\dot{w}V^{\lambda}$. The first statement of the lemma follows then immediately from the fact that if $w,w^{\prime}\in W$, and $\dot{w},\dot{w^{\prime}}\in N_{G}(T)$ are respective lifts, then, $\dot{w}\dot{w^{\prime}}$ is a lift of $ww^{\prime}$ and $\dot{w}^{-1}$ a lift of $w^{-1}$.
We prove that $W^{\lambda}\subseteq W_{\lambda}$. Let $w\in W^{\lambda}$ and $\dot{w}\in N_{G^{\lambda}}(T)$ a lift of $w$. Then, $\dot{w}G^{\lambda}\dot{w}^{-1}=G^{w\cdot\lambda}=G^{\lambda}$ and $\dot{w}V^{\lambda}=V^{w\cdot\lambda}=V^{\lambda}$ and so $w\in W_{\lambda}$. This proves the inclusion.
We now show that $W^{\lambda}$ is a normal subgroup of $W_{\lambda}$. It suffices to check that if $w\in W_{\lambda}$ and $w^{\prime}\in W^{\lambda}$, and $\dot{w}\in N_{G}(T)$, $\dot{w^{\prime}}\in N_{G^{\lambda}}(T)$ are any lifts, then $\dot{w}\dot{w^{\prime}}\dot{w}^{-1}\in G^{\lambda}$. We have $\dot{w}\dot{w^{\prime}}\dot{w}^{-1}\in\dot{w}G^{\lambda}\dot{w}^{-1}=G^{w\cdot\lambda}=G^{\lambda}$. This concludes the proof. ∎
@@
The Weyl group $W$ of $G$ acts on $\mathrm{H}^{*}(\mathrm{pt}/T)$. We just have to check that the induced action of the subgroup $W_{\lambda}$ preserves $W^{\lambda}$-invariants. This is a consequence of Lemma 4.2. ∎
@@
The group $W_{\lambda}$ acts naturally on $\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$. The subgroup $W^{\lambda}\subseteq W_{\lambda}$ acts trivially on $\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$.
@@
By definition, $G_{\lambda}$ is preserved by $W_{\lambda}$, as $G_{\lambda}$ only depends on $\overline{\lambda}\in\mathscr{P}_{V}$ and $W_{\lambda}$ is the stabilizer of $\overline{\lambda}$. The second statement follows from the fact that $G_{\lambda}$ is contained in the center of $G^{\lambda}$. ∎
We let $\overline{G^{\lambda}}\coloneqq G^{\lambda}/G_{\lambda}$. Since $G_{\lambda}$ is a normal subgroup of $G^{\lambda}$, $\overline{G^{\lambda}}$ is a reductive group. Since $G_{\lambda}$ is contained in the center of $G^{\lambda}$, the Weyl group of $\overline{G^{\lambda}}$ is isomorphic to $W^{\lambda}$.
@@
The group $W_{\lambda}$ acts naturally on $\mathrm{H}^{*}(V^{\lambda}/\overline{G^{\lambda}})$. Moreover, the $W_{\lambda}$-action on $\mathrm{H}^{*}(\mathrm{pt}/G^{\lambda})\cong\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})$ obtained in Corollary 4.3 and the $W_{\lambda}$-action on $\mathrm{H}^{*}(V^{\lambda}/\overline{G^{\lambda}})\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ obtained by the tensor product of the actions given in Lemmas 4.4 and 4.5 coincide via the isomorphism given in Proposition 3.7.
@@
We have $\mathrm{H}^{*}(V^{\lambda}/\overline{G^{\lambda}})\cong\mathrm{H}^{*}(\mathrm{pt}/(T/G_{\lambda}^{\circ}))^{W^{\lambda}}$. The restriction to $W_{\lambda}$ of the $W$ action on $T$ preserves $G_{\lambda}$ since $G_{\lambda}$ is in the center of $G^{\lambda}$. Therefore, we obtain a $W_{\lambda}$-action on $T/G_{\lambda}^{\circ}$. It induces a $W_{\lambda}$-action on $\mathrm{H}^{*}(\mathrm{pt}/(T/G_{\lambda}^{\circ}))$. The fact that $W_{\lambda}$ normalises $W^{\lambda}$ (Lemma 4.2) implies that $\mathrm{H}^{*}(\mathrm{pt}/(T/G_{\lambda}^{\circ}))^{W^{\lambda}}$ is preserved by the $W_{\lambda}$-action.
The tensor product decomposition of the action follows immediately from this and Lemma 4.4. ∎
@@
There exists a character $\varepsilon_{V,\lambda}\colon W_{\lambda}\rightarrow\mathbf{Q}^{*}$ such that for any $w\in W_{\lambda}$, $w(k_{\lambda})=\varepsilon_{V,\lambda}(w)^{-1}k_{\lambda}\in\mathrm{H}^{*}_{T}(\mathrm{pt})$.
@@
It suffices to prove the existence and unicity of the rational number $\varepsilon_{V,\lambda}(w)\in\mathbf{Q}^{*}$ satisfying the equality of the proposition, for any $w\in W_{\lambda}$. The multiplicativity then follows from the unicity of this sign.
For $w\in W_{\lambda}$, the cocharacters $\lambda,w^{-1}\cdot\lambda$ of $T$ are such that $\overline{\lambda}=\overline{w^{-1}\cdot\lambda}$. Therefore, by Lemma 2.15, there exists $\xi\in\mathbf{Q}^{*}$ such that $w(k_{\lambda})=k_{w^{-1}\cdot\lambda}=\xi k_{\lambda}$. We set $\varepsilon_{V,\lambda}(w)\coloneqq\xi^{-1}$. This proves the existence. The unicity of $\xi$ is straightforward. ∎
@@
Let $\lambda\in X_{*}(T)$, $w\in W_{\lambda}$ and $f\in\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})$. Then, $\mathrm{Ind}_{\lambda}(w\cdot f)=\varepsilon_{V,\lambda}(w)\mathrm{Ind}_{\lambda}(f)$.
@@
We now state an elementary result from the representation theory of finite groups for later use.
@@
This is [Ser71, §2.6, Théorème 8]. ∎